# Total Gluon Helicity Contribution to Proton Spin from Lattice QCD

### v1

@@ Abstract
Figure 2: The removal of ${\cal O}(a^{2}p^{2})$ discretization effect in $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$ for the ensemble C48P23. The red, green, and blue data points and their error bands represent the CDER results and their $a^{2}p^{2}\rightarrow 0$ extrapolated results using momentum modes $(0,0,p,p)$, $(0,p,p,p)$, and $(p,p,p,p)$ in (14), where $p=\big(2\pi/(aL)\big)p_{\mathrm{lat}}$ with $p_{\mathrm{lat}}$ represents momentum in the lattice unit.
@@
Numerical Results: After obtaining the renormalized matrix elements of $K^{\mu}$ and converting them to the $\overline{\rm MS}$ scheme, we are able to extract the final result by carrying out a continuum and infinite momentum extrapolation. Fig. 3 displays $\langle K^{\mu}\rangle$ as a function of lattice spacings and proton momenta. As can be seen from the plot, the results for $\mu=t,i$ exhibit a clear trend consistent with Lorentz covariance within uncertainties. Motivated by this, we adopt the following form for the continuum and infinite momentum extrapolation:
$\langle K^{\mu}\rangle=\frac{S^{\mu}}{E}\big(\Delta G+\frac{c_{\mathrm{h.t.}}}{E^{2}}\big)+c_{a}a^{2},$ (16)
@@
where $\Delta G$ is the result to be extracted, $c_{\mathrm{h.t.}}$ accounts for higher-twist contributions, and $c_{a}$ captures discretization effects. The pion mass dependence has been neglected in this analysis, primarily due to the significant computational cost of gluon matrix elements at the physical point and the very mild pion mass dependence observed in Ref. Yang:2016plb. Since the proton momenta are not significantly larger than its mass, we incorporate the higher-twist contribution using an empirical parametrization form $1/E^{2}$ in the equation above, where $E=\sqrt{M^{2}+p^{2}}$ with the proton mass $M\approx 1.02\ \mathrm{GeV}$ according to the dispersion relation. The primary sources of systematic uncertainty in our analysis are identified as follows: 1) Including or not the zero-momentum data points in the extrapolation; 2) Discretization errors associated with the finite lattice spacing; 3) The choice of the UV window in the RI/MOM scale or $a^{2}p^{2}$ (the interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ has been extended to [3.0,6.0]GeV[3.0,6.0]\ $\mathrm${GeV} to estimate the corresponding uncertainty); and 4) Systematic errors associated with the perturbative inputs (estimated by varying from three-loop to two-loop accuracy). As illustrated in FIG. 3, the joint extrapolation yields a gluon helicity contribution $\Delta G^{\overline{\mathrm{MS}},\ 10\ \mathrm{GeV}^{2}}=0.231(17)^{\mathrm{sta.}}(33)^{\mathrm{sym.}}$, which constitutes $46(7)\%$ of the proton’s total spin.
@@
Figure 3: The continuum and infinite momentum extrapolation of the total gluon helicity results using lattice spacings ranging from 0.08fm0.08\ $\mathrm${fm} to 0.11fm0.11\ $\mathrm${fm} at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$. The uncertainties are shown as separate bands: statistical uncertainty (dark gray) and all uncertainties that arise during our analysis (light gray).
@@
Summary: In this work, we presented a state-of-the-art lattice QCD calculation of the total gluon helicity contribution, $\Delta G$, to the proton spin. The calculation leveraged appropriate components of the topological current operator, $K^{\mu}$ (specifically $\mu=\{t,i\}$ where $i$ is along the direction of the proton momentum). The bare matrix element is nonperturbatively renormalized in the RI/MOM scheme and subsequently converted to the $\overline{\rm MS}$ scheme via a three-loop conversion factor with a RG improvement. The result is then extrapolated to the continuum and infinite momentum limits, yielding $\Delta G=0.231(17)^{\mathrm{sta.}}(33)^{\mathrm{sym.}}$ at $\mu^{2}=10\ \mathrm{GeV}^{2}$. This corresponds to approximately $46(7)\%$ of the proton spin and clearly rules out the possibility of a negative gluon helicity contribution Zhou:2022wzm; Whitehill:2022mpq; Hunt-Smith:2024khs. The finding offers important insight for future theoretical and experimental efforts aimed at resolving the proton spin puzzle. We have also taken into account various sources of systematic uncertainties, including those from the exclusion of zero-momentum points, discretization errors, and the choice of the renormalization scale.
@@ Acknowledgment
We thank Fangcheng He, Zhi-Cheng Hu, Xiangyu Jiang, Li-Jun Zhou, and other CLQCD and LPC members for valuable comments and suggestions. We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions Hu:2023jet, which are generated on HPC Cluster of ITP-CAS, IHEP-CAS and CSNS-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. The calculations were performed using the PyQUDA Jiang:2024lto and Chroma Edwards:2004sx software suite with QUDA Clark:2009wm; Babich:2011np; Clark:2016rdz through HIP programming model Bi:2020wpt. The numerical calculation were carried out on the ORISE Supercomputer, the Southern Nuclear Science Computing Center (SNSC). This work is supported in part by NSFC grants No. 12375080, 12525504, 12435002, 12293060, 12293062, and 12447101, No. 12205171, No. 12235008, No. 12321005, National Key R&D Program of China No.2024YFE0109800, the Ministry of Science and Technology of China under Grant No. 2024YFA1611004, by CUHK-Shenzhen under grant No. UDF01002851, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. YSBR-101, and Department of Science and Technology of Shandong province No. tsqn202312052 and 2024HWYQ-005.
Supplemental Material
.1 Detailed information of CLQCD ensembles
The results in this work, are based on the $n_{f}=2+1$ flavor ensembles from the CLQCD collaboration using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action Hu:2023jet.
The TITLS gauge action is defined in the following,
and
where $c_{\rm sw}=\frac{1}{v^{3}_{0}}$ with $v_{0}=\langle\frac{\mathrm{Re}\mathrm{Tr}\sum_{x,\mu<\nu}\mathcal{P}^{V}_{\mu\nu}(x)}{6N_{c}\tilde{V}}\rangle^{1/4}$, and $F_{\mu\nu}$ has the same form as (21). The parameters used for the simulation are described in Table 1.
.2 Bare nucleon matrix elements with topological current in the moving frame
In lattice QCD computations of the topological current,
the fundamental components requiring evaluation are the gauge potential $A_{\mu}$ and the field strength tensor $F_{\mu\nu}$, defined through the following discretized formulations:
$A_{\mu}(x+\hat{\mu}/2)=\left[\frac{U_{\mu}(x)-U_{\mu}^{\dagger}(x)}{2ig_{0}a}\right]_{\mathrm{traceless}.},$ (20)
$F_{\mu\nu}$ $=$ $\frac{i}{8g_{0}a^{2}}(\mathcal{P}_{\mu,\nu}-\mathcal{P}_{\nu,\mu}+\mathcal{P}_{\nu,-\mu}-\mathcal{P}_{-\mu,\nu}$ (21)
$+\mathcal{P}_{-\mu,-\nu}-\mathcal{P}_{-\nu,-\mu}+\mathcal{P}_{-\nu,\mu}-\mathcal{P}_{\mu,-\nu}),$
where $U_{\mu}(x)$ denotes the gauge link variable and $\mathcal{P}_{\mu,\nu}=U_{\mu}(x)U_{\nu}(x+a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x)$ represents the plaquette operator. An alternative formulation employs the full-lattice definition of the gauge potential, $\tilde{A}_{\mu}(x)=(A_{\mu}(x+\hat{\mu}/2)+A_{\mu}(x-\hat{\mu}/2))/2$. The distinction between half-lattice and full-lattice definitions manifests primarily as different discretization errors. Consequently, we adopt the half-lattice definition of the gauge potential for all subsequent calculations.
Based on the definition of local topological current in (19), one can give the expression of $K^{\mu}$ in different components, namely
$K^{i}$ $\equiv$ $K^{i}_{E\times A}+K^{i}_{A_{t}B}+K^{i}_{3A}=2\text{Tr}[A\times E]+\epsilon^{tjk}\text{Tr}[A_{t}B_{k}]-\frac{2}{3}ig_{0}\epsilon^{jkt}\mathrm{Tr}[A_{j}A_{k}A_{t}],$
$K^{t}$ $\equiv$ $K^{t}_{B\cdot A}+K^{t}_{3A}=\text{Tr}[A\cdot B]-\frac{2}{3}ig_{0}\epsilon^{ijk}\mathrm{Tr}[A_{i}A_{j}A_{k}],$ (22)
where $E_{i}=F_{ti}$ and $B_{i}=F_{jk}$. The temporal component $K^{t}$, being predominantly determined by the magnetic field strength term $K^{t}_{B\cdot A}$, exhibits matrix elements that are consistent with zero within statistical uncertainties at zero momentum, yet demonstrate a gradual increase followed by saturation to a constant value with rising momentum. Conversely, the spatial components $K^{i}$, which are primarily governed by the electric field strength term $K^{i}_{E\times A}$, maintain approximately constant matrix elements within error margins across different momentum values. The physical interpretation for this behavior aligns with (16), stemming from the treatment of protons as point-like entities: the charge density represents an intrinsic property independent of momentum, while the current density evolves from zero at rest toward a limiting value with increasing momentum. These characteristic momentum dependencies constitute a direct manifestation of Lorentz covariance within a specific reference frame.
To establish the presence of Lorentz covariance, we compute the matrix elements of the topological current operator for proton external states with varying momenta. This investigation requires two complementary computational stages: first, obtaining a stable plateau in the effective mass extraction from proton two-point correlation functions; second, calculating the disconnected three-point correlation functions with topological current insertions. Within the lattice QCD framework, appropriate smearing techniques must be employed to achieve adequate signal-to-noise ratios for both proton mass determinations and $K^{\mu}$ matrix element extractions.
FIG. 4 presents the proton effective mass behavior under ensemble C24P29. Analysis reveals that while the distillation smearing technique alone provides well-optimized plateau regions for the effective mass at low momenta, this stability deteriorates significantly at higher momenta (e.g., p=1.96GeVp=1.96\ $\mathrm${GeV}). In comparison, implementation of distillation + momentum smearing scheme yields moderately improved plateau stability at large momenta relative to the distillation-only approach.
Refer to caption Refer to caption
Figure 4: Proton effective mass with different momentum under different smearing schemes under ensemble C24P29. Left figure represents the distillation smearing scheme alone, and the right figure represents the distillation + momentum smearing. The red, green and blue points represent p=0,0.98,1.96GeVp=0,0.98,1.96\ $\mathrm${GeV}, respectively.
Employing the identical smearing methodology as proton effective mass, we extract the BMEs of $K^{t/i}$ under ensemble C24P29. Ratio data presented in FIG. 5 serve to validate the theoretical discussion surrounding (.2). In summary, the employed smearing technique demonstrates robust performance throughout the computational methodology.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Ratio $R_{K^{\mu}}(t_{f},t_{i})$ at $t_{f}\in[0.63,1.16]\ \mathrm{fm}$ and the BMEs of $K^{\mu}/(2E)$ (the gray error band) for proton external states with varying momenta under ensemble C24P29, utilizing the fitting procedure described in (9). The results are organized such that the first column corresponds to spatial components $K^{i}$, while the second column represents temporal component $K^{t}$. Rows one through three display results for proton momenta of p=0,0.98,1.96GeVp=0,0.98,1.96\ $\mathrm${GeV}, respectively.
.3 Non-perturbative renormalization with cluster decomposition error reduction (CEDR)
The computation of the renormalization constant $Z_{11}^{\mathrm{RI}}$ within the $\mathrm{RI/MOM}$ scheme, as defined in (14), necessitates the evaluation of the three-point correlation function specified in (10). This calculation presents significant numerical challenges, even with precise determinations of the gluon propagator (11) at better than $1\%$ accuracy level. The extent of this difficulty is demonstrated in FIG. 2, where the light background bands represent direct computations of $Z_{11}^{\mathrm{RI}}$ based on approximately 400 configurations of the $n_{f}=2+1$ TITLC + TITLS gauge ensemble C48P23 with lattice spacing $a\approx 0.1$ fm Hu:2023jet. These direct calculations exhibit substantial statistical uncertainties, rendering $Z_{11}^{\mathrm{RI}}$ effectively unresolved across all scales.
Applying the cluster decomposition error reduction (CDER) technique enables significant error suppression in $Z_{11}^{\mathrm{RI}}$ calculations Liu:2017man; Yang:2018bft. The fundamental principle of cluster decomposition asserts that correlators exhibit exponential decay with increasing separation between operator insertions, implying that integration beyond the correlation length primarily contributes to statistical noise rather than meaningful signal. The CDER methodology addresses this by truncating the volume integral beyond a characteristic distance, thereby achieving an enhancement of the signal-to-noise ratio by a factor of $\sqrt{V}$, as demonstrated in Liu:2017man. When implementing CDER for the computation in (10), we introduce two distinct cutoff parameters: $r_{1}$ between the gluon operator and one gauge field, and $r_{2}$ between the gauge fields within the gluon propagator. This implementation yields the modified correlator with appropriate spatial cutoffs. The CDER technique achieves computational efficiency of $\mathcal{O}(V\log V)$ through the repeated application of fast Fourier transforms (FFTs). For the three-point function specified in (10), we implement the following computational strategy:
4) Get a cutoff bare Green’s function in this way:
$G_{K}^{\mu\rho\nu,\mathrm{CDER}}(p,p)=\langle\int_{|r|<r_{1}}d^{4}r\int_{|r^{\prime}|<r_{2}}d^{4}r^{\prime}\int d^{4}xe^{ip\cdot r^{\prime}}K^{\rho}(x+r^{\prime})\mathrm{Tr}[A_{\mu}(x)A_{\nu}(x+r)]\rangle.$ (23)
Here both momentum arguments of bare Green’s function are set to $p$ in accordance with the standard implementation of the $\mathrm{RI/MOM}$ scheme, where the renormalization scale is defined by $\mu_{\rm RI}^{2}=p^{2}$.
The selection of cutoff parameters $r_{1,2}$ in the CDER technique requires careful consideration, as inappropriate choices may induce periodic oscillations in the renormalization constants Liu:2017man. FIG. 6 provides visual confirmation of this analysis. When the cutoff parameter governing the gluon field-operator correlation ($r_{1}$) is set too small (top left panel), distinct periodic oscillations become apparent in the results. These oscillations progressively stabilize as $r_{1}$ increases. A similar pattern is observed for variations in the current cutoff parameter ($r_{2}$), though changes in oscillatory behavior is less pronounced without corresponding adjustments to $r_{1}$. In FIG. 2, we employ the optimal cutoffs identified in the bottom right panel of FIG. 6, specifically $r_{1}=1.26\ \mathrm{fm}$ and $r_{2}=0.74\ \mathrm{fm}$, to perform discretization error subtraction procedure. Thus, we have successfully computed the purely gluonic Green’s function using the CDER technique, which is crucial for the non-perturbative renormalization of topological current. This method effectively suppresses the significant statistical noise generated when directly computing the three-point correlation function by truncating the volume integral beyond the characteristic distances $r_{1}$ and $r_{2}$, thus enabling the stable determination of the renormalization constant $Z_{11}^{\mathrm{RI}}$, which was previously only obtainable through the stacking statistics.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: The $\overline{\mathrm{MS}}$ renormalization constant $\bar{Z}_{11}^{\mathrm{lat}}(a^{2}p^{2},{\mu}^{2}=10\ \mathrm{GeV}^{2})$ under ensemble C48P23 as a function of $a^{2}p^{2}$ using momentum modes $(p_{\mathrm{lat}},p_{\mathrm{lat}},p_{\mathrm{lat}},p_{\mathrm{lat}})$, with different cutoffs on the gluon field-operator correlation ($r_{1}$) and current ($r_{2}$).
.4 Three-loop approximation of $\Delta G$’s off-diagonal elements
This subsection outlines the derivation of the reduced expressions presented in (15). Starting from (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD) and (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD), we obtain the exact formulations for gluon helicity in the $\overline{\mathrm{MS}}$ scheme:
$+R_{12}(a^{2}p^{2},{\mu}^{2})Z_{A}^{\mathrm{RI}}(a^{2}p^{2})\Delta\Sigma^{\mathrm{B.}}.$
In this formulation, we express the dependence on the squared renormalization scale $\mu_{\rm RI}^{2}$ of the $\mathrm{RI/MOM}$ scheme in terms of the dimensionless quantity $a^{2}p^{2}$, and the same below. We now implement a three-loop approximation for the ratio term $K^{\mathrm{lat,q}}/{J_{5}^{\mathrm{lat,q}}}$. Noting that $R_{12}$ can be determined from the topological current $K^{\mu}$ under quark external states while $R_{22}\approx 1$ is fixed by the axial vector current under quark external states, we employ the approximation:
$\big(\bar{Z}_{11}^{\mathrm{lat}}R_{12}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$ $\approx$ $\big(\bar{Z}_{11}^{\mathrm{lat}}\frac{R_{12}}{R_{22}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$ (25)
$=$ $\big(\bar{Z}_{11}^{\mathrm{lat}}\frac{K^{3\mathrm{-loop,q}}}{J_{5}^{3\mathrm{-loop,q}}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$
$\approx$ $\big(\bar{Z}_{11}^{\mathrm{lat}}\frac{K^{\mathrm{lat,q}}}{J_{5}^{\mathrm{lat,q}}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2}).$
where we used a redefined expression $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$. Furthermore, owing to the protection by chiral symmetry, the matching coefficient for the axial-vector current remains unity up to two loops, i.e., $\bar{Z}_{A}^{\mathrm{lat}}\equiv Z_{A}^{\overline{\mathrm{MS}}}\approx Z_{A}^{\mathrm{RI}}$. This convention will be consistently maintained throughout the subsequent discussion. This approximation enables the following simplified expression for gluon helicity at three-loop order:
$\Delta G^{\overline{\mathrm{MS}}}$ $\overset{3\mathrm{-loop}}{\approx}$ $\bar{Z}_{11}^{\mathrm{lat}}\big(\Delta G^{\mathrm{B.}}-R_{12}\Delta\Sigma^{\mathrm{B}.}\big)+R_{12}\bar{Z}_{A}^{\mathrm{lat}}\Delta\Sigma^{\mathrm{B.}}$ (26)
$=$ $\bar{Z}_{11}^{\mathrm{lat}}\Delta G^{\mathrm{B.}}+(\bar{Z}_{A}^{\mathrm{lat}}-\bar{Z}_{11}^{\mathrm{lat}})R_{12}\Delta\Sigma^{\mathrm{B.}}$
$\approx$ $\bar{Z}_{11}^{\mathrm{lat}}\Delta G^{\mathrm{B.}},$
where the final reduction is justified based on the hierarchy $\bar{Z}_{11}^{\mathrm{lat}}\gg(\bar{Z}_{A}^{\mathrm{lat}}-\bar{Z}_{11}^{\mathrm{lat}})R_{12}$, and by the fact that $\Delta G^{\mathrm{B.}}$ is comparable in magnitude to $\Delta\Sigma^{\mathrm{B.}}$. The derivation of expression $\Delta\Sigma^{\overline{\mathrm{MS}}}\approx\bar{Z}_{A}^{\mathrm{lat}}\Delta\Sigma^{\mathrm{B.}}$ parallels the approach used in the approximation above.
.5 Perturbative results for matching coefficients $R_{ij}$ up to three-loops
This section details the perturbative calculation of the finite renormalization (or matching) coefficients $R_{ij}$ introduced in eq. (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD) for the singlet axial-vector current $J_{5}^{\mu}$ and topological current $K^{\mu}$ from the intermediate $\mathrm{RI/MOM}$ scheme used in Lattice calculation to the $\overline{\mathrm{MS}}$ scheme typically employed in perturbative calculations in Dimensional Regularization (DR). The explicit results for $R_{ij}$ up to three loops are documented for the sake of references.

### v2

@@ Abstract
Figure 2: The removal of ${\cal O}(a^{2}p^{2})$ discretization effect in $\bar{Z}_{11}^{\mathrm{lat}}$ for the ensemble C48P23. The red, green, and blue data points and their error bands represent the CDER results and their $a^{2}p^{2}\rightarrow 0$ extrapolated results using momentum modes $(0,0,p,p)$, $(0,p,p,p)$, and $(p,p,p,p)$ in Eq. (13), where $p=\big(2\pi/(aL)\big)p_{\mathrm{lat}}$ with $p_{\mathrm{lat}}$ represents momentum in the lattice unit.
@@
Numerical Results: After obtaining the renormalized matrix elements of $K^{\mu}$ and converting them to the $\overline{\rm MS}$ scheme, we are able to extract the final result by carrying out a continuum and infinite momentum extrapolation. Fig. 3 displays $\langle K^{\mu}\rangle_{N}\equiv\bar{Z}_{11}^{\mathrm{lat.}}\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}$ as a function of lattice spacings and proton momenta. As can be seen from the plot, the results for $\mu=t,i$ exhibit a clear trend consistent with Lorentz covariance within uncertainties. Motivated by this, we adopt the following form for the continuum and infinite momentum extrapolation:
$\langle K^{\mu}\rangle_{N}=\frac{S^{\mu}}{E}\big(\Delta G+\frac{c_{\mathrm{h.t.}}}{E^{2}}\big)+c_{a}a^{2},$ (15)
@@
where $\Delta G$ is the result to be extracted, $c_{\mathrm{h.t.}}$ accounts for higher-twist contributions, and $c_{a}$ captures discretization effects. The pion mass dependence has been neglected in this analysis, primarily due to the significant computational cost of gluon matrix elements at the physical point and the very mild pion mass dependence observed in Ref. Yang et al. (2017). Since the proton momenta are not significantly larger than its mass, we incorporate the higher-twist contribution using an empirical parametrization form $1/E^{2}$ in the equation above, where $E=\sqrt{M^{2}+p^{2}}$ with the proton mass $M\approx 1\ \mathrm{GeV}$ according to the dispersion relation. The primary sources of systematic uncertainty in our analysis are identified as follows: 1) Including or not the zero-momentum data points in the extrapolation; 2) Discretization errors associated with the finite lattice spacing (the central value difference between the zero momentum $\langle K_{i}\rangle_{N}$ for the finest lattice spacing and that extrapolated to the continuum limit); 3) The choice of the UV window in the RI/MOM scale or $a^{2}p^{2}$ (the interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ has been extended to [3.0,6.0]GeV[3.0,6.0]\ $\mathrm${GeV} to estimate the corresponding uncertainty); and 4) Systematic errors associated with the perturbative input (estimated by varying from three-loop to two-loop accuracy). As illustrated in FIG. 3, the joint extrapolation yields a gluon helicity contribution $\Delta G^{\overline{\mathrm{MS}},\ 10\ \mathrm{GeV}^{2}}=0.231(17)^{\mathrm{sta.}}(44)^{\mathrm{sym.}}$, which constitutes $46(9)\%$ of the proton’s total spin.
@@
Figure 3: The continuum and infinite momentum extrapolation of the total gluon helicity is performed using lattice spacings from 0.08fm0.08\ $\mathrm${fm} to 0.11fm0.11\ $\mathrm${fm} at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$. The uncertainties are shown as separate bands: statistical uncertainty (dark gray) and all uncertainties that arise during our analysis (light gray).
@@
Summary: In this work, we presented a state-of-the-art lattice QCD calculation of the total gluon helicity contribution, $\Delta G$, to the proton spin. The calculation leveraged appropriate components of the topological current operator, $K^{\mu}$ (specifically $\mu=\{t,i\}$ where $i$ is along the direction of the proton momentum). The bare matrix element is nonperturbatively renormalized in the RI/MOM scheme and subsequently converted to the $\overline{\rm MS}$ scheme via a three-loop conversion factor with a RG improvement. The result is then extrapolated to the continuum and infinite momentum limits, yielding $\Delta G=0.231(17)^{\mathrm{sta.}}(44)^{\mathrm{sym.}}$ at $\mu^{2}=10\ \mathrm{GeV}^{2}$. This corresponds to approximately $46(9)\%$ of the proton spin and clearly rules out the possibility of a negative gluon helicity contribution Zhou et al. (2022); Whitehill et al. (2023); Hunt-Smith et al. (2024). The finding offers important insight for future theoretical and experimental efforts aimed at resolving the proton spin puzzle. We have also taken into account various sources of systematic uncertainties, including those from the exclusion of zero-momentum points, discretization errors, and the choice of the renormalization scale.
@@ Acknowledgment
We thank Fangcheng He, Zhi-Cheng Hu, Xiangyu Jiang, Li-Jun Zhou, and other CLQCD and LPC members for valuable comments and suggestions. We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions Hu and others (2024), which are generated on HPC Cluster of ITP-CAS, IHEP-CAS and CSNS-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. The calculations were performed using the PyQUDA Jiang et al. (2024) and Chroma Edwards and Joo (2005) software suite with QUDA Clark et al. (2010); Babich et al. (2011); Clark et al. (2016) through HIP programming model Bi et al. (2020). The numerical calculation were carried out on the ORISE Supercomputer, the Southern Nuclear Science Computing Center (SNSC). This work is supported in part by NSFC grants No. 12375080, 12525504, 12435002, 12293060, 12293062, and 12447101, No. 12205171, No. 12235008, No. 12321005, National Key R&D Program of China No.2024YFE0109800, the Ministry of Science and Technology of China under Grant No. 2024YFA1611004, by CUHK-Shenzhen under grant No. UDF01002851, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. YSBR-101, and Department of Science and Technology of Shandong province No. tsqn202312052 and 2024HWYQ-005.
Appendix A Supplemental Material
A.1 A. Perturbative results for matching coefficients $R_{ij}$ up to three-loops
According to Ref. Pang et al. (2024), converting the singlet axial-vector current $J_{5}^{\mu}$ and the topological current $K^{\mu}$ from the lattice $\mathrm{RI/MOM}$ scheme to the continuum $\overline{\mathrm{MS}}$ scheme requires a finite renormalization (matching) coefficient $R_{ij}$. This section details the perturbative calculation of $R_{ij}$ in dimensional regularization. Extending the results of Ref. Pang et al. (2024), we provide explicit expressions for $R_{ij}$ up to three-loop order. Note that our result for the one-loop perturbative matching coefficient differs from that reported in Ref. Pang et al. (2024), with the differences implied in the following discussion. Furthermore, we demonstrate that the off-diagonal elements of the (matching) coefficient can be neglected, leading to a simplified matching relation.
Following Ref. Pang et al. (2024), the following conditions apply within the RI/MOM renormalization scheme, namely
$\begin{pmatrix}K^{\mathrm{tree},g/q}\\ J_{5}^{\mathrm{tree},g/q}\end{pmatrix}=\begin{pmatrix}Z_{11}^{\mathrm{RI}}&Z_{12}^{\mathrm{RI}}\\ Z_{21}^{\mathrm{RI}}&Z_{22}^{\mathrm{RI}}\end{pmatrix}\begin{pmatrix}K^{\mathrm{lat.},g/q}\\ J_{5}^{\mathrm{lat.},g/q}\end{pmatrix}$, (16)
where we introduce the abbreviations $\mathcal{O}^{\mathrm{tree/lat.},{g/q}}\equiv\langle{g/q}|\mathcal{O}|{g/q}\rangle^{\mathrm{tree/lat.}}$ to simplify the notation. $|g\rangle$ or $|q\rangle$ denotes either gluon or quark external states with specific momentum and polarization. Note that $K^{\mathrm{tree},q}=J_{5}^{\mathrm{tree},g}=0$. Moreover, since only disconnected quark loops are involved, we have $J_{5}^{\mathrm{lat.},g}\approx 0$. Thus, the RI/MOM renormalization constants can be determined as:
where $\mu_{\rm RI}$ denotes the renormalization scale in the RI/MOM scheme. With the above approximation, $Z_{22}^{\mathrm{RI}}$ reduces to the axial-vector current renormalization constant $Z_{A}^{\mathrm{RI}}$ in Ref. Hu and others (2024).